Given x and y as species whose interaction is governed by

x

0 = x (( 20 0 x + 2y)

y

0 = y (( 50 + x x y),

(i) Identify the type of interaction represented by this system.

(ii) Determine the equilibrium points of this model and state the possible outcomes of

this interaction.

(iii) Linearise the system around each equilibrium point and hence discuss the nature of

the stability of each equilibrium point.

(iv) Sketch the phase portrait of the above system.


1
Expert's answer
2022-01-12T18:23:50-0500

Solution:

1)

x' and y' does not depend explicitly on time, it is autonomous system

also, this is nonlinear system


2)

"x(-20-x+2y)=0"

"y(-50+x-y)=0"


equilibrium points:

"(0,0),(0,-50),(-20,0),(120,70)"


Jacobian matrix:

"J=\begin{pmatrix}
 \partial f/ \partial x& \partial f/ \partial y \\
 \partial g/ \partial x & \partial g/ \partial y
\end{pmatrix}=\begin{pmatrix}
 -20-2x+2y & 2x \\
 y& -50+x-2y
\end{pmatrix}"


where

"f(x,y)=x(-20-x+2y)"

"g(x,y)=y(-50+x-y)"


for equilibrium points:

"J(0,0)=\begin{pmatrix}
 -20 & 0 \\
 0 & -50
\end{pmatrix}"


"\begin{pmatrix}
 x' \\
 y' 
\end{pmatrix}=\begin{pmatrix}
 -20 & 0 \\
 0 & -50
\end{pmatrix}\begin{pmatrix}
 x \\
 y
\end{pmatrix}=\begin{pmatrix}
 -20x \\
 -50y
\end{pmatrix}"


"J(0,-50)=\begin{pmatrix}
 -120 & 0 \\
 -50 & 50
\end{pmatrix}"


"\begin{pmatrix}
 x' \\
 y' 
\end{pmatrix}=\begin{pmatrix}
 -120 & 0 \\
 -50 & 50
\end{pmatrix}\begin{pmatrix}
 x \\
 y
\end{pmatrix}=\begin{pmatrix}
 -120x \\
 -50x+50y
\end{pmatrix}"


"J(-20,0)=\begin{pmatrix}
 20 & -40 \\
 0 & -70
\end{pmatrix}"


"\begin{pmatrix}
 x' \\
 y' 
\end{pmatrix}=\begin{pmatrix}
 20 & -40 \\
 0 & -70
\end{pmatrix}\begin{pmatrix}
 x \\
 y
\end{pmatrix}=\begin{pmatrix}
 20x-40y \\
 -70y
\end{pmatrix}"


"J(120,70)=\begin{pmatrix}
 -120 & 240 \\
 70 & -70
\end{pmatrix}"


"\begin{pmatrix}
 x' \\
 y' 
\end{pmatrix}=\begin{pmatrix}
 -120 & 240 \\
 70 & -70
\end{pmatrix}\begin{pmatrix}
 x \\
 y
\end{pmatrix}=\begin{pmatrix}
 -120x+240y \\
 70x-70y
\end{pmatrix}"


3)

find eigenvalues:


for "J(0,0)" :

"(-20-r)(-50-r)=0"

"r_1=-20,r_2=-50"

eigenvalues are real, r2 < r2 < 0

so, (0,0) is an asymptotically stable node


for "J(0,-50)" :

"(-120-r)(50-r)=0"

"r_1=-120,r_2=50"

eigenvalues are real, r1 < 0 < r2

so, (0,-50) is a saddle point


for "J(-20,0)" :

"(20-r)(-70-r)=0"

"r_1=20,r_2=-70"

eigenvalues are real, r2 < 0 < r1

so, (0,-50) is a saddle point


for "J(120,70)" :

"(-120-r)(-70-r)-240\cdot70=0"

"r^2+190r-8400=0"

"r=\frac{-190\pm \sqrt{190^2+4\cdot8400}}{2}"

"r_1=37,r_2=-227"

eigenvalues are real, r2 < 0 < r1

so, (120,70) is a saddle point


4)